Jackie deposited Php 75 000 in two separate accounts-a certain amount at 3% annual interest and the remaining amount at 2.5%. If at the end of the year, the interest gained from both accounts was 2 175, how much did she invest at each rate?
Answer by ikleyn(52781)

.
Jackie deposited Php 75 000 in two separate accounts-a certain amount at 3% annual interest
and the remaining amount at 2.5%. If at the end of the year, the interest gained from both accounts
was 2 175, how much did she invest at each rate?
~~~~~~~~~~~~~~~~~~~

Let x be the amount invested at 2.5% annual interest.

Then the amount invested at 3% is (75000-x).


Write an equation for the total annual interest

    0.03x + 0.025(75000-x) = 2175.    (1)


Simplify and find x

    0.03x + 1875 -0.025x = 2175

    0.03x - 0.025x = 2175 - 1875

         0.005x = 300

              x = 300/0.005 = 60,000.


Thus Php 60,000 invested at 3%, and the rest, 75,000-60,000 = 15,000 Php invested at 2.5%.   ANSWER

You may check that this answer is correct, by substituting the found values into equation (1).